OpenStudy (anonymous):

help me solve this radical: 3= 1/√x

OpenStudy (anonymous):

\[\frac{1}{\sqrt{x}}=3\] \[1=3\sqrt{x}\] \[\frac{1}{3}=\sqrt{x}\] \[(\frac{1}{3})^2=x\] \[x=\frac{1}{9}\]

OpenStudy (anonymous):

thank u